# Predicting wide binaries and deviations from standard gravity using machine learning algorithms

### v2

@@ Abstract
We present a machine learning (ML) framework for the detection of wide binary star systems and the identification of deviations from standard gravity using Gaia DR3 data. By training supervised ML models on established wide binary catalogues, we efficiently classify wide binaries and employ clustering and nearest neighbour search to pair candidate systems. We recast the problem of identifying systematic deviations from standard gravity as a supervised anomaly detection. Our approach incorporates data preprocessing techniques such as SMOTE, correlation analysis, and PCA, and achieves high accuracy and recall in both binary classification and anomaly detection tasks. The resulting publicly available code enables rapid, scalable, and customizable analysis of wide binaries and gravitational anomalies, complementing conventional analyses and providing a valuable resource for future astrophysical studies.
I Introduction and motivation
@@
Machine Learning (ML) has evolved into one of the most pivotal tools in the era of data intensive astronomy due to its efficiency and scalability, and is set to play a key role in the search for new physics in the coming decades. In recent literature, various studies have employed machine learning techniques to extract information from raw data which is otherwise difficult to analyse analytically and often computationally expensive. Stellar classification on the SIMBAD database was studied in Ref. [1], classification of accretion states of black holes was studied in Ref. [2], imposing constraints on the deviations from general relativity using ML in Ref. [3], the detection and parameter estimation process of gravitational waves using ML was carried out in Ref. [4]. For an exhaustive summary of recent works in this directions, see Refs. [5, 6, 7].
@@
In this paper, we take first steps to introduce machine learning assisted search for new physics in recently released Gaia DR3 dataset through primarily two sets of problems: (i) classification of wide binaries, and (ii) predicting deviations in the wide binaries. The former is among the well-known classification problems in astronomy, and extensively studied in literature using both traditional statistical methods [8, 9, 10] and machine learning techniques [2, 1, 7] in different contexts. The classification problem of wide binaries is interesting because these are gravitationally bound pairs of stars with large separations and can be used to study stellar evolution, dynamics, galactic structure as well as potential signatures of deviations from standard gravity [11]. Wide binary pairs of stars separated by thousands to tens of thousands of astronomical units operate precisely in the low-acceleration regime where modified gravity effects might emerge. Recent Gaia data releases have provided an unprecedented opportunity to study these systems across the Galaxy with high precision. However, identifying true gravitationally bound pairs and detecting subtle anomalies in their dynamics is complicated by noise, contamination and the scale of the dataset, thereby necessitating complex statistical analysis [8].
The latter is a relatively new problem and is rooted in the fact that despite the success of general relativity, there exist unresolved discrepancies at the galactic and cosmological scales invoking an increasing interest in modified theories of gravity, such as Modified Newtonian Dynamics (MOND) [12, 13, 14, 15, 16]. These theories predict measurable deviations from Newtonian gravity at extremely low accelerations, around $a_{o}\sim 10^{-10}m/s^{2}$ [11, 9, 17]. Wide binaries in the low acceleration regime are particularly sensitive to these deviations, as their dynamics can be significantly affected by modifications to standard gravity [18, 19, 11]. In fact, Chae [20, 21] has recently shown that a class of wide binary systems in the low-acceleration regime do exhibit systematic deviations from Newtonian gravity that is inconsistent with both dark matter and MOND based models. However, the methods for detecting these deviations often rely on probabilistic models and computationally intensive Markov Chain Monte-Carlo simulations, which struggle to scale with the increasing volume and complexity of astronomical data. Due to the relatively small number of deviating systems in an otherwise large dataset, we pose the problem of detecting deviations from gravity as the problem of anomaly detection in machine learning.
Anomaly detection technique in ML involves the detection of outliers or data points that show unexpected behaviour from the rest of the data. The conventional unsupervised anomaly detection assumes no labelled data; it relies on the assumption that anomalies are statistically different and rare compared to the majority of data. While more flexible, unsupervised methods may be less precise, especially in cases where small systematic deviations not clearly distinguishable from normal data appear [22]. Therefore, a supervised anomaly detection approach is adopted in this work because the dataset is very sparse, i.e., there are very few WBS systems in the raw data and there are very few deviations in the WBS catalogue, and it is proven in the literature [23] that unsupervised learning approaches do not perform satisfactorily with sparse datasets and require a labelled dataset approach to guide the model. In supervised anomaly detection, the model is trained on labelled data that includes both normal and anomalous instances, allowing it to learn clear distinctions between the two. This approach typically achieves higher accuracy when labelled anomalies are available, but can struggle in scenarios with rare or unseen anomalies.
@@
We have employed a supervised ML approach to predict wide binaries and potential deviations from standard gravity. As is standard in any machine learning framework, various data preprocessing techniques like correlation analysis, Synthetic Minority Oversampling Technique (SMOTE) and Principal Component Analysis (PCA) have been employed in this work. Confusion matrices and standard ML metrics have been used to analyse the performance of the models and tune the hyperparameters accordingly. We have also used the K-fold cross-validation technique for the anomaly detection problem. The codes used in this work are made available as a set of publicly available tools (hosted at https://github.com/DespCAP/G-ML) which can be used to generate a catalogue of wide binaries using our pre-trained models, and to get a rough estimate of anomalies without invoking the use of computationally expensive many-body Monte-Carlo simulations.
@@
The structure of this paper is as follows. Sec. II outlines the essential machine learning tools and techniques including that of data preprocessing and evluation of models. Sec. III describes the methodologies for using ML algorithms to predict wide binary pairs from the Gaia DR3 dataset. In Sec. IV we deal with the problem of detecting deviations from standard gravity as an anomaly detection problem. We conclude with few remarks and future outlook in Sec. V.
II Machine Learning Techniques
@@
The models are fitted onto a training dataset. The dataset utilized for prediction is called the testing dataset. The dataset was split into test and train cases in the ratio of 20: 80. The various Machine Learning models used are:
@@
    Logistic Regression: In machine learning, the Supervised Learning subcategory includes the commonly used algorithm of logistic regression. Its primary purpose is to predict the outcome of a dependent variable that belongs to a category based on a set of independent variables. This implies that the output must be categorical or discontinuous, such as Yes or No, 0 or 1, or true or false [24]. Nevertheless, rather than offering a precise value of 0 or 1, logistic regression generates probability values that fall within the range of 0 to 1.
@@
    Decision Tree Classifier: A supervised ML algorithm that is utilized primarily for classification tasks, although it can also solve regression problems. It operates on a tree-like structure that includes internal nodes representing the characteristics of a given data-set [25]. The branches denote the decision-making processes, and the leaf-nodes indicate the result. A DT comprises of two types of nodes: Decision Nodes, that possess multiple branches and are responsible for decision-making, and Leaf Nodes, which lack branches and represent the final decision or output. The Decision Tree arrives at its decisions or tests based on the properties or characteristics of the provided dataset.[25]
@@
    Random Forest Classifier: A supervised ML algorithm, helps classify the output variable as categorical or discontinuous. An RFC is based on ensemble learning, combining multiple decision trees to make more accurate predictions.[26] The algorithm creates a forest of DTs, each using a random subset (RSS). Each RSS has different features and data points. During the training process, the RFC randomly selects a subset of features, and the RSS creates a decision tree. This process is repeated several times to create multiple decision trees. [26] The algorithm predicts by aggregating each tree’s predictions and choosing the class that receives the most votes. This approach helps to improve the accuracy and robustness of the model, as it reduces the impact of individual trees that may be overfitting the data. An RFC is often used for the classification of images or text.
@@
    K-Nearest Neighbors: An ML algorithm used for the purposes of classification and regression. It falls under the category of supervised learning, meaning that it requires labelled data to train the model. It identifies new data-points on the basis of proximity to the k-nearest data-points in the training dataset [27]. The user determines the value of ‘k’ and ascertains the number of neighbours to consider. KNN is a non-parametric algorithm, meaning it makes no assumptions about the data distribution [27]. It is also easy to understand and implement, making it a popular choice for many classification and regression tasks. However, its performance can be affected by the choice of k, and it can be computationally expensive for large datasets.
@@
    Support Vector Machine: An ML algorithm, helps in tasks related to classification, regression, as well as outlier detection. It is a supervised learning algorithm; it thus requires labelled data to train the model. In SVM, the algorithm constructs a hyperplane (HP) in a High Dimensional Space (HDS) that may be deployed to separate the different classes in the data. [28] The objective is to ascertain the HP that maximises the margin, which is defined as the distance between the HP and the nearest data points of each class. Using the kernel trick (K-T) technique, SVM can handle both non-linearly and otherwise (linear) separable data (N-/LSD). The K-T transforms the input data into an HDS that can be LSD. SVM is particularly useful when dealing with high-dimensional data, for example, classifying images or text. It helps in handling datasets with a small number of samples, as it is less prone to overfitting compared to other algorithms. However, SVM may not be advisable for larger datasets and can be subject to the specific kernel function as well as other hyperparameters.[28] Nonetheless, with careful tuning of the parameters, SVM can be a powerful tool for solving many classification, regression, and outlier detection problems.
II.1 Evaluation Metrics
The accuracy, recall and F1 measure are evaluated on each ML algorithm. Confusion Matrices for all the algorithms were also plotted.
Accuracy is a metric used to determine the frequency with which a model accurately predicts the outcome of a given task. It is indicated as the ratio of the correct predictions versus the overall predictions [29]. It is particularly useful when the classes in the data are evenly distributed.
Recall is a measure of how well the model identifies positive instances. To compute this metric, the sum of true positives is divided by that of true positives and false negatives [29]. It is a valuable metric for correctly identifying all positive instances, such as in medical diagnosis.
F1 measure is a combination of precision and recall, which provides a balance between these two metrics. This metric is determined by calculating the harmonic mean of precision and recall [29]. It considers false positives as well as false negatives. F1 measure is often deployed in binary classification problems when the data is imbalanced.
@@ II.1 Evaluation Metrics
A confusion matrix is a tabular representation that is utilized to assess the effectiveness of a classification model (CM). It is a matrix that summarises the predicted and actual classifications of a model’s output, providing a more detailed view of its performance than just a single accuracy score. A confusion matrix comprises four primary components that are utilised to assess the performance of a classification model. These components are true positives (TPs), false positives (FPs), true negatives (TNs), and false negatives (FNs). Each of them carries a specific meaning. TPs signify the number of instances where the CM accurately predicts the positive-class. FPs are the number of instances where the CM predicts the positive-class despite the actual class being negative. TNs signify the number of instances where the CM accurately predicts the negative-class. FNs correspond to the number of instances where the CM predicts the negative-class despite the actual class being positive [30].
II.2 Data Preprocessing
@@
In data analysis and machine learning, having an imbalanced dataset can significantly impact the accuracy of the resulting predictions. In such cases, Synthetic Minority Oversampling Technique (SMOTE) is a commonly used method to balance the dataset. SMOTE is a technique that generates synthetic data points for the minority class to balance the distribution of the classes in the dataset. This technique creates new observations for the minority class by using interpolation methods to create ”synthetic” samples that are similar to the existing minority class observations [31]. This process continues until the minority class has a representation similar to that of the majority class. By using SMOTE to balance the dataset, the resulting distribution of the classes is more even, which allows for more accurate predictions by machine learning models. This process can mitigate the issue of imbalanced classes and can lead to better results when working with imbalanced datasets. SMOTE, thereby, helps improve the accuracy of ML models when working with imbalanced datasets, and it is frequently used in data analysis and machine learning projects.
@@
All the ML models were trained on the SMOTE-balanced as well as the raw-filtered dataset. The ML models, once trained on the SMOTE-balanced dataset, were tested on the SMOTE-balanced test dataset and the raw-filtered test dataset. As depicted in FIG. 1, there was a significant increase in accuracy and other performance metrics corresponding to each ML model. The reason for the marked increase is that the raw-filtered dataset contains very sparse entries, and the desired classes have a stark distinction. Therefore, during the training process, the ML models inherently develop a bias towards the class with a higher occurring frequency and introduce redundancies that have to be countered by training the ML models on the class-balanced dataset using the SMOTE technique [31].
Correlation analysis has been employed to quantify the degree of linear association between two continuous variables. The Pearson correlation coefficient [32], denoted as $r$, has commonly been used for this purpose. It is defined as:
$r=\frac{\sum_{i=1}^{n}(x_{i}-\bar{x})(y_{i}-\bar{y})}{\sqrt{\sum_{i=1}^{n}(x_{i}-\bar{x})^{2}}\sqrt{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}},$
where $x_{i}$ and $y_{i}$ represent individual data points, and $\bar{x}$ and $\bar{y}$ denote their respective means. The coefficient $r$ ranges from $-1$ to $1$, indicating perfect negative correlation, no correlation, and perfect positive correlation, respectively.
Prior to computing the correlation coefficient, data sets were inspected for normality and linearity, as the Pearson metric assumes both. In cases where these assumptions were violated, the Spearman rank correlation coefficient [33], a non-parametric alternative, was used instead. This approach relies on ranked data and measures monotonic relationships, regardless of linearity.
Significance of the correlation has been assessed through hypothesis testing [34], with the null hypothesis assuming no correlation between the variables ($r=0$). A two-tailed p-value has been calculated to determine whether the observed correlation differs significantly from zero, given the sample size.
Interpretation of correlation results has been guided by standard thresholds: values of $|r|<0.3$ have been considered weak, $0.70.3\leq|r|<$ 0.7 moderate, and $|r|\geq 0.7$ strong [32].
@@
II.3 Clustering and Nearest Neighbour Search
Once the set of all predicted WBS was obtained through the ML models, clustering was performed to reduce the sample size for the Nearest Neighbour Search. The K-Means clustering technique was used. K-means clustering partitions a dataset into $k$ distinct, non-overlapping clusters by minimising the within-cluster sum of squares [35]. The algorithm initializes with $k$ centroids and iteratively refines their positions by alternating between two steps: assignment of each point to the nearest centroid, and recalculation of centroid positions as the mean of all assigned points. The process converges when centroid positions stabilise or a maximum number of iterations is reached. K-means assumes clusters are spherical and approximately equal in size, which makes it sensitive to outliers and poorly suited for handling non-convex geometries or clusters with variable density. Moreover, the choice of $k$ must be specified a priori, often guided by heuristics such as the elbow method or silhouette score [35].

### v3

# Detecting wide binaries using machine learning algorithms

@@ Abstract
We present a machine learning (ML) framework for the detection of wide binary star systems using Gaia DR3 data. By training supervised ML models on established wide binary catalogues, we efficiently classify wide binaries and employ clustering and nearest neighbour search to pair candidate systems. Our approach incorporates data preprocessing techniques such as SMOTE, correlation analysis, and PCA, and achieves high accuracy and recall in the task of wide binary classification. The resulting publicly available code enables rapid, scalable, and customizable analysis of wide binaries, complementing conventional analyses and providing a valuable resource for future astrophysical studies.
keywords:
machine learning, astronomical data analysis, wide binary stars, Gaia DR3, supervised learning
1 Introduction
@@
Machine Learning (ML) has evolved into one of the most pivotal tools in the era of data intensive astronomy due to its efficiency and scalability, and is set to play a key role in the search for new physics in the coming decades. In recent literature, various studies have employed machine learning techniques to extract information from raw data which is otherwise difficult to analyse analytically and often computationally expensive. Stellar classification on the SIMBAD database was studied in Ref. [Cody et al.(2024)Cody, Scher, McDonald, Zijlstra, Alexander, and Cox], classification of accretion states of black holes was studied in Ref. [Sreehari and Nandi(2021)], imposing constraints on the deviations from general relativity using ML in Ref. [Alestas et al.(2022)Alestas, Kazantzidis, and Nesseris], the detection and parameter estimation process of gravitational waves using ML was carried out in Ref. [Koloniari et al.(2025)Koloniari, Koursoumpa, Nousi, Lampropoulos, Passalis, Tefas, and Stergioulas]. For an exhaustive summary of recent works in this directions, see Refs. [Baron(2019), Haghighi(2023), Li et al.(2025)Li, Lu, Wang, and Wang].
@@
In this paper, we take the first steps to introduce machine learning assisted search for new physics in the recently released Gaia DR3 dataset through the classification of wide binaries. The wide binary classification problem is among the well-known classification problems in astronomy, and extensively studied in literature using both traditional statistical methods [El-Badry et al.(2021)El-Badry, Rix, and Heintz, Banik and Zhao(2018), Chae(2023a)] and machine learning techniques [Sreehari and Nandi(2021), Cody et al.(2024)Cody, Scher, McDonald, Zijlstra, Alexander, and Cox, Li et al.(2025)Li, Lu, Wang, and Wang] in different contexts. The classification problem of wide binaries is interesting because these are gravitationally bound pairs of stars with large separations and can be used to study stellar evolution, dynamics, galactic structure, as well as potential signatures of deviations from standard gravity [Hernandez et al.(2012)Hernandez, Jiménez, and Allen]. Wide binary pairs of stars separated by thousands to tens of thousands of astronomical units operate precisely in the low-acceleration regime where modified gravity effects might emerge. Recent Gaia data releases have provided an unprecedented opportunity to study these systems across the Galaxy with high precision. However, identifying true gravitationally bound pairs and detecting subtle anomalies in their dynamics is complicated by noise, contamination and the scale of the dataset, thereby necessitating complex statistical analysis [El-Badry et al.(2021)El-Badry, Rix, and Heintz].
@@
We have employed a supervised ML approach to predict wide binaries. As is standard in any machine learning framework, various data preprocessing techniques like correlation analysis, Synthetic Minority Oversampling Technique (SMOTE) and Principal Component Analysis (PCA) have been employed in this work. Confusion matrices and standard ML metrics have been used to analyse the performance of the models and tune the hyperparameters accordingly. The codes used in this work are made available as a set of publicly available tools (hosted at https://github.com/DespCAP/G-ML) which can be used to generate a catalogue of wide binaries using our pre-trained models, or to train the models locally.
@@
The structure of this paper is as follows. Sec. 2 outlines the essential machine learning tools and techniques, including those of data preprocessing and evaluation of models. Sec. 3 describes the methodologies for using ML algorithms to predict wide binary pairs from the Gaia DR3 dataset. We conclude with a few remarks and a future outlook in Sec. 4.
2 Machine Learning Techniques
@@
The models are fitted onto a training dataset. The dataset utilised for prediction is called the testing dataset. The dataset was split into train and test cases in the ratio of 80:20 (train:test). The various Machine Learning models used are:
@@
    Logistic Regression: In machine learning, the Supervised Learning subcategory includes the commonly used algorithm of logistic regression. Its primary purpose is to predict the outcome of a dependent variable that belongs to a category based on a set of independent variables. This implies that the output must be categorical or discontinuous, such as Yes or No, 0 or 1, or true or false [Cox(2018)]. Nevertheless, rather than offering a precise value of 0 or 1, logistic regression generates probability values that fall within the range of 0 to 1.
@@
    Decision Tree Classifier: A supervised ML algorithm that is utilized primarily for classification tasks, although it can also solve regression problems. It operates on a tree-like structure that includes internal nodes representing the characteristics of a given data-set [Rokach(2016)]. The branches denote the decision-making processes, and the leaf-nodes indicate the result. A DT comprises of two types of nodes: Decision Nodes, that possess multiple branches and are responsible for decision-making, and Leaf Nodes, which lack branches and represent the final decision or output. The Decision Tree arrives at its decisions or tests based on the properties or characteristics of the provided dataset.[Rokach(2016)]
@@
    Random Forest Classifier: A supervised ML algorithm, helps classify the output variable as categorical or discontinuous. An RFC is based on ensemble learning, combining multiple decision trees to make more accurate predictions.[Breiman(2001)] The algorithm creates a forest of DTs, each using a random subset (RSS). Each RSS has different features and data points. During the training process, the RFC randomly selects a subset of features, and the RSS creates a decision tree. This process is repeated several times to create multiple decision trees. [Breiman(2001)] The algorithm predicts by aggregating each tree’s predictions and choosing the class that receives the most votes. This approach helps to improve the accuracy and robustness of the model, as it reduces the impact of individual trees that may be overfitting the data. An RFC is often used for the classification of images or text.
@@
    K-Nearest Neighbors: An ML algorithm used for the purposes of classification and regression. It falls under the category of supervised learning, meaning that it requires labelled data to train the model. It identifies new data-points on the basis of proximity to the k-nearest data-points in the training dataset [Cover and Hart(1967)]. The user determines the value of ‘k’ and ascertains the number of neighbours to consider. KNN is a non-parametric algorithm, meaning it makes no assumptions about the data distribution [Cover and Hart(1967)]. It is also easy to understand and implement, making it a popular choice for many classification and regression tasks. However, its performance can be affected by the choice of k, and it can be computationally expensive for large datasets.
@@
    Support Vector Machine: An ML algorithm, helps in tasks related to classification, regression, as well as outlier detection. It is a supervised learning algorithm; it thus requires labelled data to train the model. In SVM, the algorithm constructs a hyperplane (HP) in a High Dimensional Space (HDS) that may be deployed to separate the different classes in the data. [Cortes and Vapnik(1995)] The objective is to ascertain the HP that maximises the margin, which is defined as the distance between the HP and the nearest data points of each class. Using the kernel trick (K-T) technique, SVM can handle both non-linearly and otherwise (linear) separable data (N-/LSD). The K-T transforms the input data into an HDS that can be LSD. SVM is particularly useful when dealing with high-dimensional data, for example, classifying images or text. It helps in handling datasets with a small number of samples, as it is less prone to overfitting compared to other algorithms. However, SVM may not be advisable for larger datasets and can be subject to the specific kernel function as well as other hyperparameters.[Cortes and Vapnik(1995)] Nonetheless, with careful tuning of the parameters, SVM can be a powerful tool for solving many classification, regression, and outlier detection problems.
2.1 Evaluation Metrics
The accuracy, recall and F1 measure are evaluated on each ML algorithm. Confusion Matrices for all the algorithms were also plotted. Accuracy is a metric used to determine the frequency with which a model accurately predicts the outcome of a given task. It is indicated as the ratio of the correct predictions versus the overall predictions [Sokolova and Lapalme(2009)]. It is particularly useful when the classes in the data are evenly distributed. Recall is a measure of how well the model identifies positive instances. To compute this metric, the sum of true positives is divided by that of true positives and false negatives [Sokolova and Lapalme(2009)]. It is a valuable metric for correctly identifying all positive instances, such as in medical diagnosis. F1 measure is a combination of precision and recall, which provides a balance between these two metrics. This metric is determined by calculating the harmonic mean of precision and recall [Sokolova and Lapalme(2009)]. It considers false positives as well as false negatives. F1 measure is often deployed in binary classification problems when the data is imbalanced.
@@ 2.1 Evaluation Metrics
A confusion matrix is a tabular representation that is utilized to assess the effectiveness of a classification model (CM). It is a matrix that summarises the predicted and actual classifications of a model’s output, providing a more detailed view of its performance than just a single accuracy score. A confusion matrix comprises four primary components that are utilised to assess the performance of a classification model. These components are true positives (TPs), false positives (FPs), true negatives (TNs), and false negatives (FNs). Each of them carries a specific meaning. TPs signify the number of instances where the CM accurately predicts the positive-class. FPs are the number of instances where the CM predicts the positive-class despite the actual class being negative. TNs signify the number of instances where the CM accurately predicts the negative-class. FNs correspond to the number of instances where the CM predicts the negative-class despite the actual class being positive [Fawcett(2006)].
2.2 Data Preprocessing
@@
In data analysis and machine learning, having an imbalanced dataset can significantly impact the accuracy of the resulting predictions. In such cases, Synthetic Minority Oversampling Technique (SMOTE) is a commonly used method to balance the dataset. SMOTE is a technique that generates synthetic data points for the minority class to balance the distribution of the classes in the dataset. This technique creates new observations for the minority class by using interpolation methods to create ”synthetic” samples that are similar to the existing minority class observations [Chawla et al.(2002)Chawla, Bowyer, Hall, and Kegelmeyer]. This process continues until the minority class has a representation similar to that of the majority class. By using SMOTE to balance the dataset, the resulting distribution of the classes is more even, which allows for more accurate predictions by machine learning models. This process can mitigate the issue of imbalanced classes and can lead to better results when working with imbalanced datasets. SMOTE, thereby, helps improve the accuracy of ML models when working with imbalanced datasets, and it is frequently used in data analysis and machine learning projects.
@@
All the ML models were trained on the SMOTE-balanced as well as the raw-filtered dataset. The ML models, once trained on the SMOTE-balanced dataset, were tested on the SMOTE-balanced test dataset and the raw-filtered test dataset. As depicted in FIG. 1, there was a significant increase in accuracy and other performance metrics corresponding to each ML model. The reason for the marked increase is that the raw-filtered dataset contains very sparse entries, and the desired classes have a stark distinction. Therefore, during the training process, the ML models inherently develop a bias towards the class with a higher occurring frequency and introduce redundancies that have to be countered by training the ML models on the class-balanced dataset using the SMOTE technique [Chawla et al.(2002)Chawla, Bowyer, Hall, and Kegelmeyer].
Correlation analysis has been employed to quantify the degree of linear association between two continuous variables. The Pearson correlation coefficient [Benesty et al.(2009)Benesty, Chen, Huang, and Cohen], denoted as $r$, has commonly been used for this purpose. It is defined as:
$r=\frac{\sum_{i=1}^{n}(x_{i}-\bar{x})(y_{i}-\bar{y})}{\sqrt{\sum_{i=1}^{n}(x_{i}-\bar{x})^{2}}\sqrt{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}},$ (1)
where $x_{i}$ and $y_{i}$ represent individual data points, and $\bar{x}$ and $\bar{y}$ denote their respective means. The coefficient $r$ ranges from 1-1 to 0 to $1$, indicating perfect negative correlation, no correlation, and perfect positive correlation respectively.
Prior to computing the correlation coefficient, data sets were inspected for normality and linearity, as the Pearson metric assumes both. In cases where these assumptions were violated, the Spearman rank correlation coefficient [de Winter et al.(2016)de Winter, Gosling, and Potter], a non-parametric alternative, was used instead. This approach relies on ranked data and measures monotonic relationships, regardless of linearity.
Significance of the correlation has been assessed through hypothesis testing [Rainio et al.(2024)Rainio, Teuho, and Klén], with the null hypothesis assuming no correlation between the variables ($r=0$). A two-tailed p-value has been calculated to determine whether the observed correlation differs significantly from zero, given the sample size.
Interpretation of correlation results has been guided by standard thresholds: values of $|r|<0.3$ have been considered weak, $0.3\leq|r|<0.7$ moderate, and $|r|\geq 0.7$ strong [Benesty et al.(2009)Benesty, Chen, Huang, and Cohen].
@@
2.3 Clustering and Nearest Neighbour Search
Once the set of all predicted WBS was obtained through the ML models, clustering was performed to reduce the sample size for the Nearest Neighbour Search. The K-Means clustering technique was used. K-means clustering partitions a dataset into $k$ distinct, non-overlapping clusters by minimising the within-cluster sum of squares [Kanungo et al.(2002)Kanungo, Mount, Netanyahu, Piatko, Silverman, and Wu]. The algorithm initializes with $k$ centroids and iteratively refines their positions by alternating between two steps: assignment of each point to the nearest centroid, and recalculation of centroid positions as the mean of all assigned points. The process converges when centroid positions stabilise or a maximum number of iterations is reached. K-means assumes clusters are spherical and approximately equal in size, which makes it sensitive to outliers and poorly suited for handling non-convex geometries or clusters with variable density. Moreover, the choice of $k$ must be specified a priori, often guided by heuristics such as the elbow method or silhouette score [Kanungo et al.(2002)Kanungo, Mount, Netanyahu, Piatko, Silverman, and Wu].